OpenStudy (cntrl_alt_defeated):

Write an equation of a line perpendicular to y = x - 5 in slope-intercept form that passes through the point (2, -7). 1) y= x - 7 2) y= 2x + 2 3) y= -2x - 5 4) y= -2x - 3

OpenStudy (anonymous):

Note that the slope of the line y = x - 5 is 1 so we want a line whose slope is -1, and none of the options satisifies that.

OpenStudy (cntrl_alt_defeated):

so what do i do? @Ahaanomegas

OpenStudy (anonymous):

None of the options are correct. Are you sure you copied them correctly? The correct answer should be y = -x - 5.

OpenStudy (cntrl_alt_defeated):

@Ahaanomegas i left out 1/2

OpenStudy (anonymous):

So what is the new question?

OpenStudy (cntrl_alt_defeated):

it is suppossed to be 1/2x -5

OpenStudy (cntrl_alt_defeated):

but its the same question and options

OpenStudy (anonymous):

Ah. In that case, the desired slope of the line is -2, so the equation is of the form y = -2x + b, where b is a constant. To find b, note that (2, -7) is a point so \[ -7 = -2 \cdot (2) + b. \]Can you finish from here?

OpenStudy (cntrl_alt_defeated):

yes thank you very much

OpenStudy (anonymous):

You're welcome! The "best response" button is another way of saying thank you. :)